### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used for educational support, yet their response quality varies depending on the language of interaction. This paper presents an automated multilingual pipeline for generating, solving, and evaluating math problems aligned with the German K–10 curriculum. We generated 628 math exercises and translated them into English, German, and Arabic. Three commercial LLMs (GPT-4o-mini, Gemini 2.5 Flash, and Qwen-plus) were prompted to produce step-by-step solutions in each language. A held-out panel of LLM judges, including Claude 3.5 Haiku, evaluated solution quality using a comparative framework. Results show a consistent gap, with English solutions consistently rated highest, and Arabic often ranked lower. These findings highlight persistent linguistic bias and the need for more equitable multilingual AI systems in education.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly embedded in educational tools, they are reshaping how students access academic support [1, 2, 3]. In subjects like mathematics, these models are often used to explain concepts, walk through problem-solving steps, and supplement classroom instruction [1]. Nonetheless, most commercial LLMs are trained on overwhelmingly English-centric data, raising concerns about how effectively they perform in other languages [4, 5].
In multilingual educational settings, this imbalance can have serious implications. Students interacting with the same model in different languages may receive responses that vary not only in fluency, but in clarity, accuracy, and pedagogical value [1, 6]. This variation is especially problematic in math education, where precise explanations and terminology are essential. Prior studies have highlighted significant language-related disparities in LLM outputs, particularly in educational contexts [6, 5]. These findings underscore the importance of systematic, scalable evaluation across languages.
In this paper, we present a scalable, automated pipeline for evaluating how language impacts the quality of LLM-generated math solutions. Covering generation, translation, solution, and evaluation, the framework supports consistent multilingual comparison across hundreds of exercises. Using three commercial models (GPT-4o-mini (OpenAI) [7], Gemini-2.5-flash (Google) [8], and Qwen-plus (Alibaba Cloud) [9]) we apply it to English, German, and Arabic to uncover persistent performance disparities in educational contexts. This work aims to drive future research toward reducing linguistic biases in AI education tools that serve learners regardless of language background.
@@ 2 LLMs in education: Use Cases & Biases
LLMs are increasingly used in education to support both teachers and students. For educators, they assist with lesson planning, content creation, and assessment, helping reduce workload and improve efficiency [1, 2]. For students, they offer real-time, adaptive support that fosters independent learning, clarifies difficult concepts, and guides problem-solving [3].
Nonetheless, AI fairness in education is an increasing concern [10]. Growing evidence shows that LLMs introduce linguistic biases, particularly disadvantaging non-Western and low-resource language contexts [5]. Models often default to Western entities (i.e. names, foods, locations…etc) even when operating in non-Western languages like Arabic [11]. Recent work also suggests that multilingual models often perform internal reasoning in English regardless of the prompt language [4].
These tendencies raise concerns about the inclusivity and cultural relevance of AI-generated educational content. In mathematical reasoning tasks, LLMs have shown tendency to perform well in high-resource languages such as English and Chinese but struggle with languages like Korean due to difficulties processing non-English input [6]. Recent work by [5] further demonstrate this gap by evaluating six LLMs on four educational tasks across six non-English languages. They find performance drops sharply in low-resource languages, correlating with training data representation. English prompts was also shown to outperform their translations across tasks.
Such biases have real consequences. In grading and assessment contexts, models may favor Western argument structures, writing styles, and measurement systems placing non-Western students at a disadvantage. More broadly, the dominance of English-centric models reflects global educational inequality, as advanced AI tools remain most accessible to wealthier, English-speaking regions [1].
@@ 3.1 Dataset Generation & Translation
This study uses a dataset of 628 math exercises spanning five topic areas and grade levels B through H (equivalent to Grades 2–10), based on the official German K–10 mathematics curriculum [12]. [^1] Code and data are available at https://github.com/iug-htw/solution_evaluator. Exercises were generated using ChatGPT-4o (via the ChatGPT interface) in a manual, iterative process guided by curriculum-aligned topic names and learning objectives. Rather than replicating specific examples, the curriculum served as a conceptual reference to ensure broad topic coverage and grade-level appropriateness. Each item was then manually reviewed for clarity, pedagogical relevance, and alignment with the intended mathematical concept. Unclear, repetitive, or misaligned exercises were revised or removed. The resulting dataset consists of brief, one-line prompts focused on single operations or concepts, minimizing the risk for any contextual or linguistic bias. Table 1 shows example exercises.
@@ 3.2 Solution Generation
Three LLMs were used in this study: GPT-4o-mini, Gemini 2.5 Flash, and Qwen-plus. Each was prompted to generate step-by-step solutions in English, German, and Arabic. The prompts, written in the target language, asked the model to explain how to solve the given problem rather than to provide a final answer. This setup was intended to mirror how K–10 learners might seek guided support or clarification. The resulting solutions were saved in separate CSV files for each language-model combination.
@@ 4 Evaluation
To evaluate solution quality, we adopted a hybrid evaluation framework in which LLMs acted as judges. LLM-based evaluation is gaining traction as a scalable alternative to human or code-based assessments [13]. Findings suggest that LLMs can generate assessments comparable to human judgments in domains such as open-ended story generation and adversarial attack evaluations [14]. In educational contexts, LLM judgments have shown strong correlation with faculty evaluations, particularly when using pairwise comparisons and structured rubrics [15]. Other studies have shown that, with appropriate fine-tuning and prompt design, LLMs can effectively automate course evaluations across diverse university settings [16].
@@
To ensure impartiality and prevent indirect bias from models evaluating their own outputs, we implemented a held-out judging strategy. In each round, the model under evaluation was excluded from the judging panel and replaced with a neutral model, Claude 3.5 Haiku. The remaining three judges were selected from GPT-4o-mini, Gemini-2.5-flash, Qwen-plus, and Claude.
@@
During evaluation, solutions were presented in randomized order to each judge to mitigate position bias. Position bias occurs when the order of presentation influences rankings, a phenomenon observed both in LLMs and human decision-making [17, 13].
We followed a comparative assessment methodology, which has been shown to outperform direct scoring in terms of alignment with human preferences [18, 19]. Judges were asked to rank the three solutions from 1 (best) to 3 (worst) and provide a short justification to their judgment. Rankings were based on multiple educational criteria: conceptual understanding, clarity of explanation, step-by-step reasoning, correct terminology, accuracy, recognition of common mistakes, pedagogical suitability, generalizability, and grade-level appropriateness.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used for educational support, yet their response quality varies depending on the language of interaction. This paper presents an automated multilingual pipeline for generating, solving, and evaluating math problems aligned with the German K-10 curriculum. We generated 628 math exercises and translated them into English, German, and Arabic. Three commercial LLMs (GPT-4o-mini, Gemini 2.5 Flash, and Qwen-plus) were prompted to produce step-by-step solutions in each language. A held-out panel of LLM judges, including Claude 3.5 Haiku, evaluated solution quality using a comparative framework. Results show a consistent gap, with English solutions consistently rated highest, and Arabic often ranked lower. These findings highlight persistent linguistic bias and the need for more equitable multilingual AI systems in education.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly embedded in educational tools, they are reshaping how students access academic support [Yan_2023, gan2023largelanguagemodelseducation, li2024adaptinglargelanguagemodels]. In subjects like mathematics, these models are often used to explain concepts, walk through problem-solving steps, and supplement classroom instruction [Yan_2023]. Nonetheless, most commercial LLMs are trained on overwhelmingly English-centric data, raising concerns about how effectively they perform in other languages [schut2025multilingualllmsthinkenglish, gupta2025multilingualperformancebiaseslarge].
In multilingual educational settings, this imbalance can have serious implications. Students interacting with the same model in different languages may receive responses that vary not only in fluency, but in clarity, accuracy, and pedagogical value [Yan_2023, ko2025understandsolvetranslatebridging]. This variation is especially problematic in math education, where precise explanations and terminology are essential. Prior studies have highlighted significant language-related disparities in LLM outputs, particularly in educational contexts [ko2025understandsolvetranslatebridging, gupta2025multilingualperformancebiaseslarge]. These findings underscore the importance of systematic, scalable evaluation across languages.
In this paper, we present a scalable, automated pipeline for evaluating how language impacts the quality of LLM-generated math solutions. Covering generation, translation, solution, and evaluation, the framework supports consistent multilingual comparison across hundreds of exercises. Using three commercial models (GPT-4o-mini (OpenAI) [openai2024gpt4omini], Gemini-2.5-flash (Google) [google2025gemini25flash], and Qwen-plus (Alibaba Cloud) [alibaba2025qwenplus]) we apply it to English, German, and Arabic to uncover persistent performance disparities in educational contexts. This work aims to drive future research toward reducing linguistic biases in AI education tools that serve learners regardless of language background.
@@ 2 LLMs in education: Use Cases & Biases
LLMs are increasingly used in education to support both teachers and students. For educators, they assist with lesson planning, content creation, and assessment, helping reduce workload and improve efficiency [Yan_2023, gan2023largelanguagemodelseducation]. For students, they offer real-time, adaptive support that fosters independent learning, clarifies difficult concepts, and guides problem-solving [li2024adaptinglargelanguagemodels].
Nonetheless, AI fairness in education is an increasing concern [simbeck2024]. Growing evidence shows that LLMs introduce linguistic biases, particularly disadvantaging non-Western and low-resource language contexts [gupta2025multilingualperformancebiaseslarge]. Models often default to Western entities (i.e. names, foods, locations…etc) even when operating in non-Western languages like Arabic [naous2024havingbeerprayermeasuring]. Recent work also suggests that multilingual models often perform internal reasoning in English regardless of the prompt language [schut2025multilingualllmsthinkenglish].
These tendencies raise concerns about the inclusivity and cultural relevance of AI-generated educational content. In mathematical reasoning tasks, LLMs have shown tendency to perform well in high-resource languages such as English and Chinese but struggle with languages like Korean due to difficulties processing non-English input [ko2025understandsolvetranslatebridging]. Recent work by [gupta2025multilingualperformancebiaseslarge] further demonstrate this gap by evaluating six LLMs on four educational tasks across six non-English languages. They find performance drops sharply in low-resource languages, correlating with training data representation. English prompts was also shown to outperform their translations across tasks.
Such biases have real consequences. In grading and assessment contexts, models may favor Western argument structures, writing styles, and measurement systems placing non-Western students at a disadvantage. More broadly, the dominance of English-centric models reflects global educational inequality, as advanced AI tools remain most accessible to wealthier, English-speaking regions [Yan_2023].
@@ 3.1 Dataset Generation & Translation
This study uses a dataset of 628 math exercises spanning five topic areas and grade levels B through H (equivalent to Grades 2-10), based on the official German K-10 mathematics curriculum [RahmenlehrplanMathematik2015]. [^1] Code and data are available at https://github.com/iug-htw/solution_evaluator. Exercises were generated using ChatGPT-4o (via the ChatGPT interface) in a manual, iterative process guided by curriculum-aligned topic names and learning objectives. Rather than replicating specific examples, the curriculum served as a conceptual reference to ensure broad topic coverage and grade-level appropriateness. Each item was then manually reviewed for clarity, pedagogical relevance, and alignment with the intended mathematical concept. Unclear, repetitive, or misaligned exercises were revised or removed. The resulting dataset consists of brief, one-line prompts focused on single operations or concepts, minimizing the risk for any contextual or linguistic bias. Table 1 shows example exercises.
@@ 3.2 Solution Generation
Three LLMs were used in this study: GPT-4o-mini, Gemini 2.5 Flash, and Qwen-plus. Each was prompted to generate step-by-step solutions in English, German, and Arabic. The prompts, written in the target language, asked the model to explain how to solve the given problem rather than to provide a final answer. This setup was intended to mirror how K-10 learners might seek guided support or clarification. The resulting solutions were saved in separate CSV files for each language-model combination.
@@ 4 Evaluation
To evaluate solution quality, we adopted a hybrid evaluation framework in which LLMs acted as judges. LLM-based evaluation is gaining traction as a scalable alternative to human or code-based assessments [zheng2023judgingllmasajudgemtbenchchatbot]. Findings suggest that LLMs can generate assessments comparable to human judgments in domains such as open-ended story generation and adversarial attack evaluations [chiang-lee-2023-large]. In educational contexts, LLM judgments have shown strong correlation with faculty evaluations, particularly when using pairwise comparisons and structured rubrics [ishida2024largelanguagemodelspartners]. Other studies have shown that, with appropriate fine-tuning and prompt design, LLMs can effectively automate course evaluations across diverse university settings [yuan2024explorationhighereducationcourse].
@@
To ensure impartiality and prevent indirect bias from models evaluating their own outputs, we implemented a held-out judging strategy. In each round, the model under evaluation was excluded from the judging panel and replaced with a neutral model, Claude 3.5 Haiku [anthropic2025claude35]. The remaining three judges were selected from GPT-4o-mini, Gemini-2.5-flash, Qwen-plus, and Claude.
@@
During evaluation, solutions were presented in randomized order to each judge to mitigate position bias. Position bias occurs when the order of presentation influences rankings, a phenomenon observed both in LLMs and human decision-making [tan2024judgebenchbenchmarkevaluatingllmbased, zheng2023judgingllmasajudgemtbenchchatbot].
We followed a comparative assessment methodology, which has been shown to outperform direct scoring in terms of alignment with human preferences [liu2025aligninghumanjudgementrole, liusie2024llmcomparativeassessmentzeroshot]. Judges were asked to rank the three solutions from 1 (best) to 3 (worst) and provide a short justification to their judgment. Rankings were based on multiple educational criteria: conceptual understanding, clarity of explanation, step-by-step reasoning, correct terminology, accuracy, recognition of common mistakes, pedagogical suitability, generalizability, and grade-level appropriateness.
@@ Acknowledgements.
Declaration on Generative AI
During the preparation of this work, the authors used OpenAI’s GPT-4.1 to check grammar and spelling and to enhance the writing style. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the publication’s content.